tigerlover (tigerlover):

What is the slope of the line shown below? A. 1 B. -1 C. -6 D. 6

OpenStudy (ninja4535):

M=y2-y1/x2-x1 = 9-2/3-(4) =7/7 or 1

563blackghost (563blackghost):

Agreed.

OpenStudy (3mar):

The slope of a straight line passes through two points is calculated from the relation: \[m=\frac{ y_2-y_2 }{ x_2-x_1 }=\frac{ 9-2 }{ 3-(-4) }=\frac{ 7 }{ 7 }=1\] This is obvious that the line has an angle of 45 with x-axis. Sp we can choose the first choice (A)

OpenStudy (3mar):

Thanks, Ninja4535.